OpenStudy (anonymous):

(1.07)^4x = 5/2 Solve for x.

OpenStudy (anonymous):

\[1.07^{4x}\] or \[1.07^{4}x\] ???

OpenStudy (anonymous):

^(4x)

OpenStudy (anonymous):

ok first step it to take the log of both sides log ((1.07)^(4x)) = log (5/2) what's the next step? remember: log(a)^b = b*log(a)

OpenStudy (anonymous):

slight correction remember: log(a^b) = b*log(a)

OpenStudy (jdoe0001):

\(\bf 1.07^{4x}=\cfrac{5}{2}\\ \quad \\ \textit{log cancellation of }\quad log_aa^x = x\\ \quad \\ log_{1.07}(1.07^{4x})=log_{1.07}\left(\cfrac{5}{2}\right)\implies 4x = log_{1.07}\left(\cfrac{5}{2}\right)\\ \quad \\ \textit{change of base rule}\quad log_ab=\cfrac{log_ca}{log_cb}\\ \quad \\ 4x = \cfrac{log_{10}\left(\frac{5}{2}\right)}{log_{10}(1.07)}\)

OpenStudy (jdoe0001):

\(\bf \textit{change of base rule}\quad log_ab=\cfrac{log_cb}{log_ca}\) to be exact =)